# Spectral Concentration at the Edge of Stability: Information Geometry of Kernel Associative Memory

### v2

@@ 5.2 Geometric Derivation of Antagonism
This explains the “amplification” mechanism: even a small intrinsic mismatch ($\tilde{\nabla}L$) is magnified by the huge curvature $\lambda_{1}$, resulting in a massive Euclidean gradient $\nabla L$.
@@ 6.2 Implications for Deep Learning Generalization
The phenomenon of Spectral Concentration may offer broader insights into the generalization mechanics of over-parameterized neural networks. Recent studies in deep learning have reported a “feature learning” regime, often referred to as Neural Collapse [10], where the weight spectrum becomes low-rank to capture task-relevant structure. Our results provide a geometric explanation for this: the network minimizes the effective dimensionality (via the Fisher Information spectrum) to ensure robust inference (Dual Equilibrium), effectively performing a geometric Occam’s razor.
@@ 6.3 The Geometry of Attention Mechanisms
Modern Hopfield Networks are mathematically equivalent to the attention mechanism in Transformers [11]. Our geometric framework implies that the effectiveness of attention layers may also rely on specific spectral structures in their Key-Query matrices. If attention heads also operate at an Edge of Stability to sharpen their focus, our duality theory (Euclidean force vs. Riemannian curvature) could provide a new tool for analyzing and stabilizing the training of large-scale Transformer models. Future work will investigate whether trained Transformers exhibit similar spectral signatures to those found on the Ridge.

### v3

@@ 5.2 Geometric Derivation of Antagonism
This explains the “amplification” mechanism: the huge curvature $\lambda_{1}$ magnifies small intrinsic mismatches into massive Euclidean gradients.
@@ 6.2 Implications for Deep Learning Generalization
The phenomenon of Spectral Concentration may offer broader insights into the generalization mechanics of over-parameterized neural networks. Recent studies report a “feature learning” regime, or Neural Collapse [10], where the weight spectrum becomes low-rank to capture task-relevant structure. Notably, Karakida et al. [11, 12] demonstrated that deep neural networks inherently exhibit a pathological FIM spectrum dominated by outliers. Our findings suggest that KLR learning exploits this spectral pathology to stabilize memory. Our results provide a geometric explanation for this: the network minimizes the effective dimensionality (via the Fisher Information spectrum) to ensure robust inference (Dual Equilibrium), effectively performing a geometric Occam’s razor.
@@ 6.3 The Geometry of Attention Mechanisms
Modern Hopfield Networks are mathematically equivalent to the attention mechanism in Transformers [13]. Our framework suggests that attention layers may also rely on specific spectral structures in their Key-Query matrices. If attention heads operate at an Edge of Stability, our duality theory could provide new tools for analyzing large-scale Transformers. Future work will investigate whether they exhibit spectral signatures similar to the Ridge.